# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v3

@@ 7.1 Parallelism, GPU, and JAX engineering
A major line of work accelerates KANs by making their core operations GPU-friendly. Iterative B-spline evaluations are replaced by non-iterative ReLU-power activations [50] or reformulated as CUDA-optimized matrix kernels [129, 139]. Model-level restructuring further boosts throughput: dual-matrix compression with Random Fourier Features [73], custom backward kernels to reduce memory stalls [199], and GPU-accelerated B-spline evaluation [130]. Architectural parallelism also plays a role, including multi-branch KAN–MLP hybrids [60] and domain-decomposition designs for PINNs [39, 128, 196]. At the framework level, JAX implementations [67, 124] leverage jit/XLA and optimized autodiff to deliver substantial end-to-end speedups.
@@ 8 Sparsity & Regularization
eliminating redundant connections without accuracy loss.
@@ 9.3 NTK-Based Convergence and Physics-Informed Extensions
giving a rigorous rate under appropriate learning rates. The same holds in expectation for SGD.
@@ 10 Practical “Choose–Your–KAN” Guide
When to prefer KANs over MLPs. KANs are most advantageous when the learning problem benefits from explicit control over the underlying function space. In practice, KANs are preferable to standard MLPs when (i) data are sparse or expensive and fast convergence is required, (ii) a shallow architecture is desired for interpretability or stability, or (iii) prior knowledge about the solution structure is available. Examples include globally smooth or oscillatory functions (Chebyshev/Fourier bases), localized or multi-scale features (splines or wavelets), and nonsmooth or discontinuous behavior (Sinc- or rational-based KANs). In contrast, when large amounts of data are available and no structural prior is known, generic MLPs may remain competitive due to their simplicity and hardware efficiency.
@@ 11 Current Gaps and Path Forward
Comparisons of splines, Chebyshev, Gaussians, etc. [41] highlight problem dependence but mask deeper structure. Each basis involves choices of normalization, grids, and regularization—too many degrees of freedom for a single benchmark to be conclusive.

### v4

@@ 7.1 Parallelism, GPU, and JAX engineering
A major line of work accelerates KANs by making their core operations GPU-friendly. Iterative B-spline evaluations are replaced by non-iterative ReLU-power activations [53] or reformulated as CUDA-optimized matrix kernels [132, 142]. Model-level restructuring further boosts throughput: dual-matrix compression with Random Fourier Features [76], custom backward kernels to reduce memory stalls [202], and GPU-accelerated B-spline evaluation [133]. Architectural parallelism also plays a role, including multi-branch KAN–MLP hybrids [63] and domain-decomposition designs for PINNs [42, 131, 199]. At the framework level, JAX implementations [70, 127] leverage jit/XLA and optimized autodiff to deliver substantial end-to-end speedups, while specialized hardware designs such as computing-in-memory architectures using Gaussian-like memory cells further improve energy-efficient KAN inference [298].
@@ 8 Sparsity & Regularization
eliminating redundant connections without accuracy loss. Application-driven sparsity can also arise from the data representation itself; for example, SP-KAN employs sparse sinusoidal perception layers aligned with the intrinsic sparsity of infrared targets [296] [^53] https://github.com/xdFai.
@@ 9.3 NTK-Based Convergence and Physics-Informed Extensions
giving a rigorous rate under appropriate learning rates. The same holds in expectation for SGD. Intuitively, a better-conditioned NTK (i.e., a larger and more balanced spectrum of eigenvalues) implies that gradients propagate more uniformly across training directions, leading to faster and more stable convergence in practice.
@@ 10 Practical “Choose–Your–KAN” Guide
When to prefer KANs over MLPs. It is important to note that highly optimized MLP implementations remain the preferred choice in many high-throughput production settings where inference latency and hardware efficiency are the primary concerns. KANs are most advantageous when the learning problem benefits from explicit control over the underlying function space. In practice, KANs are preferable to standard MLPs when (i) data are sparse or expensive and fast convergence is required, (ii) a shallow architecture is desired for interpretability or stability, or (iii) prior knowledge about the solution structure is available. Examples include globally smooth or oscillatory functions (Chebyshev/Fourier bases), localized or multi-scale features (splines or wavelets), and nonsmooth or discontinuous behavior (Sinc- or rational-based KANs). In contrast, when large amounts of data are available and no structural prior is known, generic MLPs may remain competitive due to their simplicity and hardware efficiency.
@@ 11 Current Gaps and Path Forward
Comparisons of splines, Chebyshev, Gaussians, etc. [44] highlight problem dependence but mask deeper structure. Each basis involves choices of normalization, grids, and regularization—too many degrees of freedom for a single benchmark to be conclusive. For fair comparisons, future studies should standardize key variables such as grid resolution, normalization schemes, training settings, and evaluation metrics.